# Bound-electron self-energy calculations in Feynman and Coulomb gauges: detailed analysis

### v1

@@ Appendix D Derivative of the free-electron self-energy operator
where the coefficients $N_{1}$, $N_{2}$ and $N_{3}$ can be obtained by differentiating the formula (52):
$\begin{gathered}N_{1}=\frac{\partial^{2}a(p_{0},p)}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}},\\ N_{2}=\varepsilon_{a}\frac{\partial^{2}b(p_{0},p)}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}+\frac{\partial b(p_{0},p)}{\partial p_{0}}\Bigr|_{p_{0}=\varepsilon_{a}}+\frac{\partial^{2}c(p_{0},p)}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}},\\ N_{3}=\frac{\partial^{2}b(p_{0},p)}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}\end{gathered}$ (79)
In the Feynman gauge, the derivatives of the coefficients $a$ and $b$ are:
$\frac{\partial^{2}a(p_{0},p)}{\partial p_{0}^{2}}=-\frac{8}{m(1-\rho)}\left\{1+\frac{1}{1-\rho}\left[\ln\rho-\frac{2p_{0}^{2}}{m^{2}}\left(\frac{1+\rho}{\rho}+\frac{2}{1-\rho}\ln\rho\right)\right]\right\},$ (80)
$\frac{\partial b(p_{0},p)}{\partial p_{0}}=\frac{2p_{0}}{m^{2}(1-\rho)^{2}}\left\{3-\rho+\frac{2}{1-\rho}\ln\rho\right\},$ (81)
$\frac{\partial^{2}b(p_{0},p)}{\partial p_{0}^{2}}=\frac{2}{m^{2}(1-\rho)^{2}}\left\{3-\rho+\frac{2}{1-\rho}\left[\ln\rho-\frac{p_{0}^{2}}{m^{2}}\left(\frac{2+5\rho-\rho^{2}}{\rho}+\frac{6}{1-\rho}\ln\rho\right)\right]\right\}.$ (82)
@@
$\begin{gathered}N_{1}=-2m\int_{0}^{1}dx\frac{\partial^{2}\ln Y}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}},\\ N_{2}=2\int_{0}^{1}dx(1-x)\left(\varepsilon_{a}\frac{\partial^{2}\ln Y}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}+2\frac{\partial\ln Y}{\partial p_{0}}\Bigr|_{p_{0}=\varepsilon_{a}}\right),\\ N_{3}=2\int_{0}^{1}dx\left((1-x)\frac{\partial^{2}\ln Y}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}-\sqrt{x}\int_{0}^{1}du\frac{\partial^{2}\ln Z}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}\right),\end{gathered}$ (84)
@@
$\begin{gathered}\frac{\partial\ln Y}{\partial p_{0}}=\frac{2p_{0}(x-1)}{m^{2}-(p_{0}^{2}-p^{2})(1-x)-i0},\\ \frac{\partial^{2}\ln Y}{\partial p_{0}^{2}}=\frac{2(x-1)(m^{2}+(p_{0}^{2}+p^{2})(1-x))}{(m^{2}-(p_{0}^{2}-p^{2})(1-x)-i0)^{2}},\\ \frac{\partial^{2}\ln Z}{\partial p_{0}^{2}}=\frac{2(u-1)(m^{2}+p_{0}^{2}(1-u)+p^{2}(1-xu))}{(m^{2}-p_{0}^{2}(1-u)+p^{2}(1-xu)-i0)^{2}}.\end{gathered}$ (85)
@@
$\begin{gathered}\int_{0}^{1}du\left(\frac{\partial^{2}\ln Z}{\partial p_{0}^{2}}\right)=2\Big[\frac{p_{0}^{4}-3p_{0}^{2}(m^{2}+p^{2})-xp^{2}(m^{2}+p^{2}-3p_{0}^{2})}{m^{2}\rho(p_{0}^{2}-xp^{2})^{2}}\\ -\frac{(m^{2}+p^{2}(1-x))(3p_{0}^{2}+xp^{2})}{(p_{0}^{2}-xp^{2})^{3}}\ln\left(\frac{m^{2}+p^{2}-p_{0}^{2}}{m^{2}+p^{2}(1-x)}\right)\Big].\end{gathered}$ (86)
@@
$\begin{gathered}\int_{0}^{1}dy\frac{y}{(m^{2}-(p_{0}^{2}-p^{2})y)^{2}}=\frac{1}{p_{0}^{2}-p^{2}}\left(\frac{\ln\rho}{p_{0}^{2}-p^{2}}+\frac{1}{m^{2}\rho}\right),\\ \int_{0}^{1}dy\frac{y^{2}}{(m^{2}-(p_{0}^{2}-p^{2})y)^{2}}=\frac{m^{2}}{(p_{0}^{2}-p^{2})^{3}}\left(\frac{1}{\rho}+2\ln\rho-\rho\right),\\ \int_{0}^{1}dy\frac{y^{3}}{(m^{2}-(p_{0}^{2}-p^{2})y)^{2}}=\frac{1}{(p_{0}^{2}-p^{2})^{4}}(p_{0}^{4}-2(p_{0}^{2}+2m^{2})p^{2}+\frac{2m^{4}}{\rho}+6m^{4}\ln\rho\\ +4m^{2}p_{0}^{2}+(p^{2})^{2}-2m^{4}),\\ \int_{0}^{1}dy\frac{y^{2}}{m^{2}-(p_{0}^{2}-p^{2})y}=-\frac{1}{2}\left(\frac{1}{p_{0}^{2}-p^{2}}+\frac{2m^{2}}{(p_{0}^{2}-p^{2})^{2}}+\frac{2m^{4}\ln\rho}{(p_{0}^{2}-p^{2})^{3}}\right).\end{gathered}$ (87)

### v2

@@ Appendix D Derivative of the free-electron self-energy operator
where the coefficients $N_{1}$, $N_{2}$, and $N_{3}$ can be obtained by differentiating the formula (56):
$\begin{gathered}N_{1}=\frac{\partial^{2}a(p_{0},p)}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}},\\ N_{2}=\varepsilon_{a}\frac{\partial^{2}b(p_{0},p)}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}+\frac{\partial b(p_{0},p)}{\partial p_{0}}\Bigr|_{p_{0}=\varepsilon_{a}}+\frac{\partial^{2}c(p_{0},p)}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}},\\ N_{3}=\frac{\partial^{2}b(p_{0},p)}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}.\end{gathered}$ (87)
In the Feynman gauge, the derivatives of the coefficients $a$ and $b$ are (see Sapirstein and Cheng (2023); Malyshev et al. (2024)):
$\frac{\partial^{2}a(p_{0},p)}{\partial p_{0}^{2}}=-\frac{8}{m_{e}(1-\rho)}\left\{1+\frac{1}{1-\rho}\left[\ln\rho-\frac{2p_{0}^{2}}{m_{e}^{2}}\left(\frac{1+\rho}{\rho}+\frac{2}{1-\rho}\ln\rho\right)\right]\right\},$ (88)
$\frac{\partial b(p_{0},p)}{\partial p_{0}}=\frac{2p_{0}}{m_{e}^{2}(1-\rho)^{2}}\left\{3-\rho+\frac{2}{1-\rho}\ln\rho\right\},$ (89)
$\frac{\partial^{2}b(p_{0},p)}{\partial p_{0}^{2}}=\frac{2}{m_{e}^{2}(1-\rho)^{2}}\left\{3-\rho+\frac{2}{1-\rho}\left[\ln\rho-\frac{p_{0}^{2}}{m_{e}^{2}}\left(\frac{2+5\rho-\rho^{2}}{\rho}+\frac{6}{1-\rho}\ln\rho\right)\right]\right\}.$ (90)
@@
$\begin{gathered}N_{1}=-2m_{e}\int_{0}^{1}dx\frac{\partial^{2}\ln Y}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}},\\ N_{2}=2\int_{0}^{1}dx(1-x)\left(\varepsilon_{a}\frac{\partial^{2}\ln Y}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}+2\frac{\partial\ln Y}{\partial p_{0}}\Bigr|_{p_{0}=\varepsilon_{a}}\right),\\ N_{3}=2\int_{0}^{1}dx\left((1-x)\frac{\partial^{2}\ln Y}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}-\sqrt{x}\int_{0}^{1}du\frac{\partial^{2}\ln Z}{\partial p_{0}^{2}}\Bigr|_{p_{0}=\varepsilon_{a}}\right),\end{gathered}$ (92)
@@
$\begin{gathered}\frac{\partial\ln Y}{\partial p_{0}}=\frac{2p_{0}(x-1)}{m_{e}^{2}-(p_{0}^{2}-p^{2})(1-x)-i0},\\ \frac{\partial^{2}\ln Y}{\partial p_{0}^{2}}=\frac{2(x-1)(m_{e}^{2}+(p_{0}^{2}+p^{2})(1-x))}{(m_{e}^{2}-(p_{0}^{2}-p^{2})(1-x)-i0)^{2}},\\ \frac{\partial^{2}\ln Z}{\partial p_{0}^{2}}=\frac{2(u-1)(m_{e}^{2}+p_{0}^{2}(1-u)+p^{2}(1-xu))}{(m_{e}^{2}-p_{0}^{2}(1-u)+p^{2}(1-xu)-i0)^{2}}.\end{gathered}$ (93)
@@
$\begin{gathered}\int_{0}^{1}du\left(\frac{\partial^{2}\ln Z}{\partial p_{0}^{2}}\right)=2\Big[\frac{p_{0}^{4}-3p_{0}^{2}(m_{e}^{2}+p^{2})-xp^{2}(m_{e}^{2}+p^{2}-3p_{0}^{2})}{m_{e}^{2}\rho(p_{0}^{2}-xp^{2})^{2}}\\ -\frac{(m_{e}^{2}+p^{2}(1-x))(3p_{0}^{2}+xp^{2})}{(p_{0}^{2}-xp^{2})^{3}}\ln\left(\frac{m_{e}^{2}+p^{2}-p_{0}^{2}}{m_{e}^{2}+p^{2}(1-x)}\right)\Big].\end{gathered}$ (94)
@@
$\begin{gathered}\int_{0}^{1}dy\frac{y}{(m_{e}^{2}-(p_{0}^{2}-p^{2})y)^{2}}=\frac{1}{p_{0}^{2}-p^{2}}\left(\frac{\ln\rho}{p_{0}^{2}-p^{2}}+\frac{1}{m_{e}^{2}\rho}\right),\\ \int_{0}^{1}dy\frac{y^{2}}{(m_{e}^{2}-(p_{0}^{2}-p^{2})y)^{2}}=\frac{m_{e}^{2}}{(p_{0}^{2}-p^{2})^{3}}\left(\frac{1}{\rho}+2\ln\rho-\rho\right),\\ \int_{0}^{1}dy\frac{y^{3}}{(m_{e}^{2}-(p_{0}^{2}-p^{2})y)^{2}}=\frac{1}{(p_{0}^{2}-p^{2})^{4}}(p_{0}^{4}-2(p_{0}^{2}+2m_{e}^{2})p^{2}+\frac{2m_{e}^{4}}{\rho}+6m_{e}^{4}\ln\rho\\ +4m_{e}^{2}p_{0}^{2}+(p^{2})^{2}-2m_{e}^{4}),\\ \int_{0}^{1}dy\frac{y^{2}}{m_{e}^{2}-(p_{0}^{2}-p^{2})y}=-\frac{1}{2}\left(\frac{1}{p_{0}^{2}-p^{2}}+\frac{2m_{e}^{2}}{(p_{0}^{2}-p^{2})^{2}}+\frac{2m_{e}^{4}\ln\rho}{(p_{0}^{2}-p^{2})^{3}}\right).\end{gathered}$ (95)